Gold which has a density of 19.32g/cm^3, is the most ductile metal and can be pressed into a thin leaf or drawn out into a long fiber.
a) if a sample of gold , with a mass of 27.63 g , is pressed into a leaf of 1.000 micrometer thickness , what is the area of the leaf ?
b)if instead, the gold is drawn out into a cylindrical fiber of radius 2.500 micrometer, What is the lenght of fiber?
1
Expert's answer
2018-09-10T12:56:08-0400
a)
m=ρV=ρhA
A=m/ρh=27.63/((19.32)(0.0001))=14300 cm^2=1.43 m^2.
b)
l=V/A=(m/ρ)/(πr^2 )
l=(27.63/((19.32)))/(π〖(0.00025)〗^2 )=7.284∙〖10〗^6 cm=72.84 km.